OpenStudy (anonymous):

Find the average rate of change of each function on the interval specified. f(x) = x^2 on [1,5]

OpenStudy (anonymous):

I know that x^2 is the shape of a parabola.

OpenStudy (anonymous):

average rate of change just means slope

OpenStudy (anonymous):

so just find the slope between (1,1) and (5,25)

OpenStudy (anonymous):

m = ysub2-ysub1/xsub2/xsub1

OpenStudy (anonymous):

how do you format that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

at the bottom of the page, there is button that says Equation. Click on it

OpenStudy (anonymous):

so I understand where you got the 1 and 5 how did you get 1,1 and 5,25?

OpenStudy (anonymous):

solving for slope is not difficult :)

OpenStudy (anonymous):

f(1) = 1^2 = 1 f(5) = 5^2 = 25 so (1,1) and (5,25)

OpenStudy (anonymous):

\[25-1/5-1 = 24/4 = 6\]

OpenStudy (anonymous):

probably a better way to format that :)

OpenStudy (anonymous):

average rate of change is 6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am still getting the hang of the format

OpenStudy (anonymous):

how do you make it look all nice?

OpenStudy (anonymous):

there are bunch of formats and symbols on the right and side when you click Equation button. Just play with it

OpenStudy (anonymous):

that helped out thanks!